# Asymptotic Analysis of High Order IMEX-RK Methods for ES-BGK Model at Navier-Stokes level

### v2

@@ 3 Asymptotic Properties of the IMEX RK Schemes
We conclude this section with the following observations. If conditions (3.73) are not satisfied, order reduction is likely to occur in these intermediate regimes of $\varepsilon$. Therefore, we adopt IMEX-RK schemes of type II that satisfy conditions (3.73) ensuring the validity of the estimates in Theorem 3.8 and preventing order reduction in the intermediate regime.
In this paper, we examine two IMEX-RK schemes of type ARS, introduced in [7] and called IMEX-II-GSA3 and IMEX-II-ISA3. These two schemes are specifically designed in [7] to improve the resolution of hyperbolic relaxation systems in the intermediate regime $\varepsilon\approx\Delta t$.
Before presenting these two schemes, we note that from the second condition in (3.73), under the assumption of $\tilde{b}=b$ we get
@@
which equals zero if $\tilde{a}_{s1}=a_{s1}$.
In Section 4, we will show that IMEX-II-GSA3 scheme, which is GSA and $\tilde{b}\neq b$ leads to a mild order reduction for small values of $\varepsilon$. On the contrary, IMEX-II-ISA3 with $\tilde{b}=b$, does not exhibit order reduction, as it satisfies (3.90), (as confirmed by the numerical results in the next section).
Remark 3.9.
Although the results of Theorem 3.7 in [21] pertain to IMEX-RK schemes of type II, the authors, in that theorem, assume the hypothesis that $\mathcal{G}(f^{(i)})=\mathcal{G}(f^{(n)})+\mathcal{O}(\Delta t)$ and, in the proof, they also assume that $\tau^{(i)}=\tau^{n}+\mathcal{O}(\Delta t)$.
In contrast, in our Theorem 3.8 we do not make these assumptions; instead, we require the additional conditions (3.73). However, if these conditions are not satisfied, the estimation of the local truncation error will include an additional term of order $\mathcal{O}(\varepsilon\Delta t)$ which also appears in the local truncation error in Theorem 3.7 in [21].
Therefore, as previously mentioned, by appropriately choosing a type II scheme, for example type ARS, with $\tilde{b}=b$, these additional conditions can be automatically satisfied, allowing us to guarantee results of Theorem 3.8.
@@ 4 Numerical Results
Throughout this section, we consider IMEX-RK schemes of type I and II. In particular, we take into account the GSA IMEX-II-GSA3 scheme and IMEX-II-ISA3 with $\tilde{b}=b$. For comparison, we consider the IMEX-RK(4,3,3) scheme of type I not GSA in [10], and the scheme GSA ARS(4,4,3) of type II in [1].
@@
  • Third-order IMEX-RK(4,3,3) scheme of type I in [10]. (Top: explicit method. Bottom: implicit method).
@@
$\begin{split}\rho_{0}(x)&=1+0.2\sin(\pi x),\quad u_{0}(x)=\left(1,0\right),\quad T_{0}(x)=\frac{1}{\rho_{0}(x)}\cr f_{0}(x,v)&=\mathcal{M}(x,v)-\frac{\varepsilon}{\tau}\left(I-\Pi_{\mathcal{M}}\right)(v_{1}\partial_{x}\mathcal{M})\end{split}$ (4.94)
@@
$\mathcal{M}(x,v)=\frac{\rho_{0}(x)}{2\pi T_{0}(x)}\exp\left(-\frac{|v-u_{0}(x)|^{2}}{2T_{0}(x)}\right)$
@@
Table 4.1 shows results for third order accurate IMEX-RK schemes at $t=0.25$. First we can see order reduction in the intermediate regime for ARS(4,4,3) and IMEX-RK(4,3,3) scheme. Although we observe order reductions for IMEX-RK at small Knudsen numbers, such as $10^{-4}$, it is a typical issue of IMEX-RK methods. In general, it is highly nontrivial to avoid order reduction of IMEX-RK schemes in the intermediate regime (for example, see papers [5, 6, 11]).
@@
$\mathcal{M}(x,v)=\frac{\rho_{0}(x)}{2\pi T_{0}(x)}\exp\left(-\frac{(v_{1}-u_{0}(x))^{2}+v_{2}^{2}}{2T_{0}(x)}\right)$ (4.95)
For the velocity discretization, we use $32\times 32$ uniform points in the domain $v\in[-10,10]\times[-10,10]$. Note that, for the IMEX-RK(4,3,3) scheme in order to achieve the expected order of convergence, we required more points in velocity. In this test we use $N_{v}=80$ uniform points for the velocity. We use $N_{x}$ uniform points for space in the interval $x\in[0,2]$ with periodic boundary conditions. The time step is taken as $\Delta t=\text{CFL}\Delta x/10$ with $\text{CFL}=0.9$, $\tau=1$ and $\nu=-1/2$.
@@
Table 4.2 shows the result for the third order accurate IMEX-RK schemes at $t=0.25$. As the one dimensional case we can see order reduction in the intermediate regime ($\varepsilon=10^{-4}$) for the ARS(4,4,3) scheme. The third order scheme IMEX-II-ISA3 achieves the third order accuracy for all the values of $\varepsilon$ even in the intermediate regime, whereas IMEX-II-GSA3 and IMEX-RK(4,3,3) schemes show a slight order reduction for the value $\varepsilon=10^{-4}$.
@@
$\left(\rho_{0},u_{x0},u_{y0},T_{0}\right)=\Biggl\{\begin{array}[]{cc}(1,M\sqrt{2},0,1),&-1\leq x\leq 0.5,\\[5.69054pt] (\frac{1}{8},0,0,\frac{1}{4}),&\textrm{otherwise},\end{array}$
@@
We consider these transport coefficients when computing the reference solutons to NSE. We remark that this choice matches the viscosity of ES-BGK model and the one derived from BTE for Maxwellian molecules [18].
In Figs. 4.1-4.2, we present numerical results for BTE, NSE and ES-BGK model for different Knudsen numbers $\varepsilon=0.5,0.1$ respectively. For BTE, we use an explicit fourth-order semi-Lagrangian scheme [10], while for NSE, we use the classical RK4 and WENO23. For ES-BGK, we use the combination of IMEX-II-ISA3 for the time discrsetization and WENO23 for the space one. For $\varepsilon=0.5$, the numerical solution of the ES-BGK model is very close to that of BTE, while solutions to NSE are deviated from the other solutions. On the other hand, for $\varepsilon=0.1$, the discrepancy between three solutions become smaller, which confirms the consistency of the three models.
Figure 4.1: Comparison between reference solutions of BTE and NSEs with the numerical solutions for ES-BGK model. We report the case of density, velocity, temperature and heat flux at time t=0.1, 0.25, 0.4t=0.1,\,0.25,\,0.4. The Knudsern number is $\varepsilon=0.5$.
@@
In Fig. 4.3, we compare the numerical solution of the ES-BGK model with those of the compressible Navier-Stokes equations and the compressible Euler equations. The reference solution of the Euler equations and of the Navier-Stokes equations was generated by using spectral method for the spacial derivatives and fourth-order RK method for the time on a grid of 3200 for $\varepsilon=0,10^{-2},10^{-4}$ respectively. As we can see from the Fig. 4.3 the numerical solution of the ES-BGK model is very close to that of the NS equations. Furthermore for a better visualization, we zoomed the two solutions on a portion of the domain.
@@ 5 Conclusions
In this work, we investigated the asymptotic behavior at the Navier-Stokes level, reproducing theoretical results, Theorems 3.6-3.8. We showed that existing IMEX-RK schemes including type I and II can accurately capture the NS limit without resolving the small scales determined by the Knudsen number. To address the issues on order reduction of these schemes at the Navier-Stokes level, we propose IMEX-RK schemes of type II developed in [7]. In particular, the IMEX-II-ISA3 scheme of type ARS satisfying additional conditions (3.73) guarantees uniform accuracy avoiding order reduction. Finally, we provided numerical examples that validate the theoretical results.
@@ A.1 Basic properties of Gaussian distribution function
A.2 Chapmann-Enskog expansion of ES-BGK model
@@
$\langle\phi\mathcal{M}(f)\rangle+\langle\phi v\cdot\nabla_{x}\mathcal{M}(f)\rangle=-\varepsilon\nabla_{x}\langle\phi vf_{1}\rangle,$
@@
$\partial_{t}U+\nabla_{x}F(U)=-\varepsilon\nabla_{x}H({\bf U})$ (A.106)
@@
$\begin{split}\tau(f_{1}-g)&=(I-\Pi_{\mathcal{M}})(\tau(f_{1}-g))\cr&=-(I-\Pi_{\mathcal{M}})(\partial_{t}\mathcal{M}+v\cdot\nabla_{x}\mathcal{M})+\mathcal{O}(\varepsilon)\cr&=-(I-\Pi_{\mathcal{M}})(v\cdot\nabla_{x}\mathcal{M})+\mathcal{O}(\varepsilon)\cr\end{split}$ (A.113)
@@
$V=\frac{v-u}{\sqrt{T}},\quad A(V)=V\otimes V-\frac{1}{d_{v}}|V|^{2}Id,\quad B(V)=\frac{1}{2}V(|V|^{2}-(d_{v}+2)),$
@@
$\left\langle v\phi g\right\rangle=\mathcal{G}(U),\quad\textrm{where}\quad\mathcal{G}(U)=(0,\nu\rho\Theta_{1},\nu\rho\Theta_{1}u)^{\top},$ (A.117)
@@
$\begin{split}\left\langle v\phi f_{1}\right\rangle&=\left\langle v\phi g\right\rangle-\frac{1}{\tau}\left\langle v\phi\mathcal{M}\left(A(V):\sigma(u)+2B(V)\cdot\nabla_{x}\sqrt{T}\right)\right\rangle+\mathcal{O}(\varepsilon),\\ &H(U)=\mathcal{G}(U)-\frac{1}{\tau}\mathcal{F}(U)+\mathcal{O}(\varepsilon),\end{split}$ (A.120)
@@
$\kappa=\frac{d_{v}+2}{2}\frac{p}{\tau}$

### v3

@@ 3 Asymptotic Properties of the IMEX RK Schemes
The same conclusions regarding the local truncation error presented in Theorem 3.6 can be applied here for the case of type II. ∎
3.3 Uniform accuracy
In the stiff case, i.e., $\Delta t\gg\varepsilon$, the phenomenon of order reduction may occur, particularly in the worst case when $\Delta t\approx\varepsilon$ (intermediate regime of $\varepsilon$). In this scenario, the classical order of the method can drop off, leading to a loss of accuracy in IMEX-RK schemes within this regime. This phenomenon was investigated in detail in [11, 8, 7]. In particular in [8] the authors considered a prototype hyperbolic relaxation system and performed an asymptotic expansion up to $\mathcal{O}(\varepsilon)$ for IMEX-RK methods of type I and II. They showed that, under additional order conditions, these schemes effectively reduce to explicit-type RK methods at the $\mathcal{O}(\varepsilon)$ level. These methods were carefully designed to accurately match the $\mathcal{O}(\varepsilon)$ terms up to a desired order $p$, ensuring that effectively behave as explicit RK schemes of order $p$ for the convection-diffusion equation when $\varepsilon$ is small but not negligible. However, constructing high-order IMEX-RK schemes of type I requires additional stages to satisfy extra-order conditions, their construction is particularly challenging, [8, 7]. In this work, we therefore focus on IMEX-RK schemes of type II,[8], since their structures make them easier to construct.
We note that if conditions (3.68) are not satisfied, order reduction is likely to occur in these intermediate regimes of $\varepsilon$, for example, in the numerical tests (Test 1 and Test 2), a classical ARS scheme of type II in (4.87) show this phenomenon of order reduction. Alternatively, we adopt IMEX-RK schemes of type II introduced in [8] that satisfy conditions (3.68) ensuring the consistency of the schemes (Theorem 3.8). Furthermore, these schemes satisfy additional order conditions, preventing order reduction in the intermediate regime and ensuring uniform accuracy over the full range of Knudsen numbers. In particular, we examine two IMEX-RK schemes of type II called IMEX-II-GSA3 (4.88) and IMEX-II-ISA3 (4.89). These two schemes are specifically designed in [8] to improve the resolution of hyperbolic relaxation systems, especially in the intermediate regime, i.e., $\varepsilon\approx\Delta t$.
In Section 4, we compare these methods with other IMEX-RK schemes of type I or II existing in literature.
Finally, we note that IMEX-II-ISA3 scheme fulfills $\tilde{b}=b$ and the quantity in (3.84) becomes
@@ 3.3 Uniform accuracy
which equals zero if $\tilde{a}_{s1}=a_{s1}$. It is worth noting that the assumption $\tilde{b}=b$, plays an important rule in mitigating the lost of accuracy, especially in regimes where $\varepsilon\approx\Delta t$, as demonstrated in [11]. In Section 4, we will show that IMEX-II-GSA3 scheme, which is GSA and $\tilde{b}\neq b$ leads to a mild order reduction for small values of $\varepsilon$. On the contrary, IMEX-II-ISA3 with $\tilde{b}=b$, does not exhibit order reduction, (as confirmed by the numerical results in the next section).
@@ 4 Numerical Results
Throughout this section, we consider IMEX-RK schemes of type I and II. In particular, we take into account the GSA IMEX-II-GSA3 scheme and IMEX-II-ISA3 with $\tilde{b}=b$. For comparison, we consider the SI-IMEX(4,4,3) scheme of type I not GSA in [7], and the scheme GSA ARS(4,4,3) of type II in [1].
@@
  • Third-order SI-IMEX(4,4,3) scheme of type I in [7]. (Top: explicit method. Bottom: implicit method).
@@
$\begin{split}\rho_{0}(x)&=1+0.2\sin(\pi x),\quad{u_{0}(x)=1},\quad T_{0}(x)=\frac{1}{\rho_{0}(x)},\cr f_{0}(x,v)&=\mathcal{M}(x,v)-\frac{\varepsilon}{\tau}\left(I-\Pi_{\mathcal{M}}\right)(v_{1}\partial_{x}\mathcal{M})\end{split}$ (4.90)
@@
$\mathcal{M}(x,v)=\frac{\rho_{0}(x)}{2\pi T_{0}(x)}\exp\left(-\frac{|v-u_{0}(x)|^{2}}{2T_{0}(x)}\right).$
@@
Table 4.1 shows results for third order accurate IMEX-RK schemes at $t=0.25$. First we can see order reduction in the intermediate regime for ARS(4,4,3) and SI-IMEX(4,4,3) scheme. Although we observe order reductions for IMEX-RK at small Knudsen numbers, such as $10^{-4}$, it is a typical issue of IMEX-RK methods. In general, it is highly nontrivial to avoid order reduction of IMEX-RK schemes in the intermediate regime (for example, see papers [11, 12, 9]).
@@
$\mathcal{M}(x,v)=\frac{\rho_{0}(x)}{2\pi T_{0}(x)}\exp\left(-\frac{(v_{1}-u_{0}(x))^{2}+v_{2}^{2}}{2T_{0}(x)}\right).$ (4.91)
For the velocity discretization, we use $32\times 32$ uniform points in the domain $v\in[-10,10]\times[-10,10]$. Note that, for the SI-IMEX(4,4,3) scheme in order to achieve the expected order of convergence, we required more points in velocity. In this test we use $N_{v}=80$ uniform points for the velocity. We use $N_{x}$ uniform points for space in the interval $x\in[0,2]$ with periodic boundary conditions. The time step is taken as $\Delta t=\text{CFL}\Delta x/10$ with $\text{CFL}=0.9$, $\tau=1$ and $\nu=-1/2$.
@@
Table 4.2 shows the result for the third order accurate IMEX-RK schemes at $t=0.25$. As the one dimensional case we can see order reduction in the intermediate regime ($\varepsilon=10^{-4}$) for the ARS(4,4,3) scheme. The third order scheme IMEX-II-ISA3 achieves the third order accuracy for all the values of $\varepsilon$ even in the intermediate regime, whereas IMEX-II-GSA3 and SI-IMEX(4,4,3) schemes show a slight order reduction for the value $\varepsilon=10^{-4}$.
At the theoretical level, the ES-BGK model approximates the NSE equations with an error of order $\mathcal{O}(\varepsilon^{2})$, as is well known from the Chapman-Enskog expansion. However, verifying this convergence numerically is extremely challenging. Indeed, the numerical method approximates the ES-BGK model, not the NSE system directly. Hence, the observed error will always be a combination of the modeling error, between ES-BGK and NSE, and the numerical discretization error in approximating both models. With this view point, to show the error between ES-BGK and NSE is $\mathcal{O}(\varepsilon^{2})$, we consider the same initial data used in Test 2, compute very accurate solutions for each model for a range of Knudsen numbers, and check whether the discrepancy of the numerical solutions show expected order $\mathcal{O}(\varepsilon^{2})$. In view of this, we compute numerical solutions with sufficiently small $\Delta t$ and $\Delta x$ (for both ES-BGK and NSE) and large number of velocity grid points (for ES-BGK), and measure the discrepancy in relative $L^{1}$-norm. In the following figure 4.1, it appears that the convergence rates are close to $2$ when $10^{-4}\leq\varepsilon\leq 10^{-2}$. On the other cases, the second order is not obtained. Note that when $\varepsilon$ is relatively small $(\varepsilon<10^{-4})$, the result is improved as we increase the accuracy of numerical solutions. However, when $\varepsilon$ is relatively large $(\varepsilon>10^{-2})$, the errors are not improved even for high accurate solutions. This can be interpreted as modelling errors between ES-BGK and NSE.
Figure 4.1: Knudsen number($\varepsilon$) vs relative $L_{1}$-norm. For the comparison, we used the same initial data for test 2, and compute numerical solutions up to $t=0.5$.
@@
$\left(\rho_{0},u_{x0},u_{y0},T_{0}\right)=\begin{cases}(1,M\sqrt{2},0,1),&-1\leq x\leq 0.5,\\[5.69054pt] (\frac{1}{8},0,0,\frac{1}{4}),&\hskip 11.38109pt\textrm{otherwise},\end{cases}$
@@
We consider these transport coefficients when computing the reference solutions to NSE. We remark that this choice matches the viscosity of ES-BGK model and the one derived from BTE for Maxwellian molecules [18].
In Figs. 4.2-4.3, we present numerical results for BTE, NSE and ES-BGK model for different Knudsen numbers $\varepsilon=0.5,0.1$ respectively. For BTE, we use an explicit fourth-order semi-Lagrangian scheme [5], while for NSE, we use the classical RK4 and WENO23. For ES-BGK, we use the combination of IMEX-II-ISA3 for the time discretization and WENO23 for the space one. For $\varepsilon=0.5$, the numerical solution of the ES-BGK model is very close to that of BTE, while solutions to NSE are deviated from the other solutions. On the other hand, for $\varepsilon=0.1$, the discrepancy between three solutions become smaller, which confirms the consistency of the three models.
Figure 4.2: Comparison between reference solutions of BTE and NSEs with the numerical solutions for ES-BGK model. We report the case of density, velocity, temperature and heat flux at time t=0.1, 0.25, 0.4t=0.1,\,0.25,\,0.4. The Knudsen number is $\varepsilon=0.5$.
@@
In Fig. 4.4, we compare the numerical solution of the ES-BGK model with those of the compressible Navier-Stokes equations and the compressible Euler equations. The reference solution of the Euler equations and of the Navier-Stokes equations was generated by using spectral method for the spatial derivatives and fourth-order RK method for the time on a grid of 3200 for $\varepsilon=0,10^{-2},10^{-4}$ respectively. As we can see from the Fig. 4.4 the numerical solution of the ES-BGK model is very close to that of the NS equations. Furthermore for a better visualization, we zoomed the two solutions on a portion of the domain.
@@ 5 Conclusions
In this work, we investigated the asymptotic behavior at the Navier-Stokes level, reproducing theoretical results, Theorems 3.6 and Theorem 3.8. We showed that existing IMEX-RK schemes including type I and II can accurately capture the NS limit without resolving the small scales determined by the Knudsen number. To address the issues on order reduction of these schemes at the Navier-Stokes level, we propose IMEX-RK schemes of type II developed in [8]. In particular, the IMEX-II-ISA3 scheme of type ARS satisfying additional conditions (3.68) guarantees uniform accuracy avoiding order reduction. Finally, we provided numerical examples that validate the theoretical results.
@@ A.1 Basic properties of Gaussian distribution function
A.2 Chapman-Enskog expansion of ES-BGK model
@@
$\langle\phi\mathcal{M}(f)\rangle+\langle\phi v\cdot\nabla_{x}\mathcal{M}(f)\rangle=-\varepsilon\nabla_{x}\cdot\langle\phi vf_{1}\rangle,$
@@
$\partial_{t}U+\nabla_{x}F(U)=-\varepsilon\nabla_{x}\cdot H({\bf U})$ (A.102)
@@
$\begin{split}\tau(f_{1}-g)&=(I-\Pi_{\mathcal{M}})(\tau(f_{1}-g))\cr&=-(I-\Pi_{\mathcal{M}})(\partial_{t}\mathcal{M}+v\cdot\nabla_{x}\mathcal{M})+\mathcal{O}(\varepsilon)\cr&=-(I-\Pi_{\mathcal{M}})(v\cdot\nabla_{x}\mathcal{M})+\mathcal{O}(\varepsilon).\end{split}$ (A.109)
@@
$V=\frac{v-u}{\sqrt{T}},\quad A(V)=V\otimes V-\frac{1}{d_{v}}|V|^{2}Id,\quad B(V)=\frac{1}{2}V(|V|^{2}-(d_{v}+2)),$ (A.111)
@@
$\left\langle v\phi g\right\rangle={G}(U),\quad\textrm{where}\quad{G}(U)=(0,\nu\rho\Theta_{1},\nu\rho\Theta_{1}u)^{\top},$ (A.114)
@@
$\begin{split}\left\langle v\phi f_{1}\right\rangle&=\left\langle v\phi g\right\rangle-\frac{1}{\tau}\left\langle v\phi\mathcal{M}\left(A(V):\sigma(u)+2B(V)\cdot\nabla_{x}\sqrt{T}\right)\right\rangle+\mathcal{O}(\varepsilon),\\ &H(U)={G}(U)-\frac{1}{\tau}\mathcal{F}(U)+\mathcal{O}(\varepsilon),\end{split}$ (A.117)
@@
$\kappa=\frac{d_{v}+2}{2}\frac{p}{\tau},$